Fermi National Laboratory
 

Accelerator Update for October 13, 2006 - October 16, 2006

 

Friday October 13

The day shift began with the Tevatron (TeV) and Antiproton Source (Pbar) off due to magnet replacements, with the Recycler, Linac, Booster, and the Main Injector (MI) off due to an access in Linac, with experts accessing the TeV F sector, with MiniBooNE off due to Booster being off, and with NuMI off due to problems with HV101A. The Collider Detector at Fermilab (CDF) and D-Zero (D0) are both off and in access.

Operations prepared the Booster for an access at 9:27 AM. Technicians will make vacuum repairs.

Power supply experts completed their repair of HV101A. NuMI got its first pulse of beam at 12:05 PM.

The Booster access ended and MiniBooNE resumed taking beam at 12:30 PM.

The Neutron Therapy Facility (NTF) reported at 12:34 PM that they had finished treating their last patient for the day.

The two pictures below show the interface between the two TeV superconducting magnet, D32-5 dipole and the D33-1 quadrupole. You can see the superconductor bus in both of these images; the picture on the right uses a glove to highlight the failed bus.

Both the quadrupole and dipole need to be replaced. Techs will also replace D35-1A, which is a spool piece that has a dead dipole trim element.

MI experts began conducting slip-injection studies around 3 PM.

Power supply experts completed their Recycler work. A Recycler expert took some protons at 3:49 PM to take measure beam lifetime.

Controls experts replaced a card for NuMI power supply HV101A at 4:47PM. NuMI was turned off for this change.

NuMI resumed taking beam at 6:54 PM, but then they discovered a noise problem that had to be tracked down, but the problem wasn’t found. Investigations continued.

The Cryo Coordinator reported at 9:53 PM that the cryo system temperature at TeV sector D1 warmed up to liquid nitrogen level.

Saturday October 14

The midnight shift began with the TeV and Pbar in access, and with NuMI and MiniBooNE taking beam.

At 1:23 AM, NuMI tripped off for the first of two times in four hours due to HV101 surpassing its upper current limit. It reset each time.

At 3:32 AM, a Booster power supply (WAPS) tripped off for the first of three times in fifteen minutes. It reset okay.

At 6:55 AM, WAPS tripped off for the first of three times in twelve minutes. It reset okay.

At 8:58 AM, the upper VCB for MI-60 opened on a transformer oil alarm. Operations called in experts. The experts found a digital meter that had failed. They power cycled the meter. Beam to NuMI resumed at 10:26 AM.

TeV techs reported at 11:12 AM that they had replaced the D35-1 spool piece.

A Preaccelerator expert switched to the H- Source at 12:21 PM.

At 3:58 PM, Pbar experts reported that they found a pipe with too many welds and would not hold a vacuum. They will have to begin preparing a new vacuum pipe tonight for installation on Sunday morning.

TeV techs reported at 6:03 PM that they had completed their installation of the D32-5 dipole magnet and had began their installation of the D33-1 quadrupole. They found carbon residue in the beam pipe. TeV dipoles 32-2 through 4 have to be moved for further investigations. TeV experts called in an alignment crew.

Sunday October 15

The midnight shift began with the TeV and Pbar in access, and with NuMI and MiniBooNE taking beam.

At 12:19 AM, WAPS tripped off for the first of five times in about 2.5 hours. It reset each time.

At 12:22 AM, TeV techs reported that they had to remove D33-1 to install a vacuum pump.

At 3:02 AM, Operations reported that the NuMI and MiniBooNE beam permits had been tripping off due an Interface Rack Monitor (IRM) for a Booster RF station (BRF13). Operations called in an expert, who arrived at 4:04 AM, but then BRF13 tripped. The expert worked on the IRM timing card.

At 6:13 AM, Operations reported the BRF13 IRM communications had resumed, but then a damper problem showed up. Operations investigated and with the help of an expert discovered a bad damper amplifier. The amplifier was replaced and beam returned to NuMI and MiniBooNE at 6:41 AM.

At 6:49 AM, the BRF13 IRM went down again. Operations called in an expert.

Operations reported at 8:29 AM that the clock input for the BRF13 IRM was changed and the station was back up.

Operations reported that the damper amplifier had been replaced and that beam through Booster had resumed at 8:52 AM.

At 10:52 AM, the Duty Mechanic reported a lot of makeup flow for the Pbar LCW system. Operations found no leaks in the areas operators could search without halting beam to NuMI and MiniBooNE. Operations contacted experts.

At 11:38 AM, a Booster RF station (BRF18) tripped off and wouldn’t reset. An expert came in and worked on the station.

At 12:28 PM, experts accessed Pbar’s AP2 and 3 enclosure (these two beamlines share the same enclosure) and found an inch of water on the floor. The leak was traced to a faulty hose gasket on a quadrupole magnet (IQ5). The access and leak repair halted beam to NuMI and MiniBooNE due to the closeness of the Pbar Transport enclosure (AP2&3) and the MI-8 line transport line, which runs between Booster and MI.

Technicians repaired the IQ5 LCW leak. NuMI and MiniBooNE resumed taking beam at 1:56 PM.

At 3:44 PM, WAPS tripped off for the first of three times in 15 minutes. It reset each time.

Operations ran beam to the Pbar dump at 6:56 PM to verify the ability to stack.

At 8:41 PM, WAPS tripped off twice in ten minutes. It reset both times. Operations contacted an expert.

Monday October 16

The midnight shift began with the TeV in access and with NuMI and MiniBooNE taking beam.

At 5:49 AM, BRF17 tripped off. It reset okay.

The Plans for Monday

  1. Replace the Tevatron D3 magnets
  2. Check the Tevatron separator vacuum at sectors D0 and D17 – the TeV separators may require a bake
  3. Complete safety system tests

Accelerator Update Archive

More Information
For Tevatron luminosity charts and the current status of Fermilab's accelerators and detectors (live!), please go to Fermilab Now.

Comments and Suggestions
What do you think about the Accelerator Updates? Please send comments and suggestions to: accelupdates@fnal.gov.

last modified 10/16/2006    email Fermilab